# ExeCoder: Empowering Large Language Models with Executability Representation for Code Translation

### v2

@@ Appendix C Example in XLCoST-Instruct
Figure 10. Example for learning code semantic in XLCoST-Instruct.
@@
Figure 11. Example for learning functional semantic in XLCoST-Instruct.
Prompt: Translate the given code from python to cpp. The input contains the source code and a Abstract Syntax Tree of the code. The input Code is marked with <Code>and </Code>. Please note that the code entered is a complete program with main fuction. The Abstract Syntax Tree of the code is marked with <AST>and </AST>.
@@
Input: <Code> def Kmultiples ( n , k ) : a = n for i in range ( 1 , k + 1 ) : print ( "{}*{}={}".format ( n , i , a ) ) j = 0 while ( n >= ( 1 << j ) ) : a += n & ( 1 << j ) j += 1 </Code> <AST> G describes a graph among nodes 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, and 115. In this graph: Node 5 represents code def. Node 6 represents code Kmultiples. Node 8 represents code :. Node 13 represents code (. Node 14 represents code n. Node 15 represents code ,. Node 16 represents code k. Node 17 represents code ). Node 20 represents code N. Node 21 represents code =. Node 22 represents code 16. Node 23 represents code K. Node 24 represents code =. Node 25 represents code 7. Node 26 represents code Kmultiples. Node 29 represents code for. Node 30 represents code i. Node 31 represents code in. Node 33 represents code :. Node 35 represents code (. Node 36 represents code N. Node 37 represents code ,. Node 38 represents code K. Node 39 represents code ). Node 40 represents code a. Node 41 represents code =. Node 42 represents code n. Node 43 represents code range. Node 48 represents code (. Node 49 represents code 1. Node 50 represents code ,. Node 52 represents code ). Node 55 represents code while. Node 57 represents code :. Node 59 represents code k. Node 60 represents code +. Node 61 represents code 1. Node 62 represents code print. Node 64 represents code j. Node 65 represents code =. Node 66 represents code 0. Node 67 represents code (. Node 69 represents code ). Node 72 represents code (. Node 74 represents code ). Node 75 represents code n. Node 76 represents code >=. Node 82 represents code (. Node 84 represents code ). Node 85 represents code a. Node 86 represents code +=. Node 88 represents code j. Node 89 represents code +=. Node 90 represents code 1. Node 92 represents code .. Node 93 represents code format. Node 94 represents code (. Node 95 represents code n. Node 96 represents code ,. Node 97 represents code i. Node 98 represents code ,. Node 99 represents code a. Node 100 represents code ). Node 101 represents code 1. Node 102 represents code <<. Node 103 represents code j. Node 104 represents code n. Node 105 represents code &. Node 107 represents code ". Node108representscode{}*{}={}.Node109representscode". Node 110 represents code (. Node 112 represents code ). Node 113 represents code 1. Node 114 represents code <<. Node 115 represents code j. In this graph: Node 0 is connected to nodes 1, 2, 3, 4. Node 1 is connected to nodes 2, 3, 4, 5, 6. Node 2 is connected to node 3. Node 3 is connected to node 4. Node 4 is connected to node 5. Node 7 is connected to nodes 8, 9, 10, 11, 12. Node 9 is connected to nodes 10, 11. Node 10 is connected to nodes 11, 12, 13. Node 11 is connected to nodes 12, 13, 14. Node 12 is connected to nodes 13, 14. Node 18 is connected to node 19. Node 19 is connected to nodes 20, 21, 22, 23, 24, 25. Node 27 is connected to nodes 28, 29, 30, 31, 32. Node 28 is connected to nodes 29, 30, 31. Node 32 is connected to nodes 33, 34. Node 34 is connected to nodes 35, 36, 37. Node 44 is connected to nodes 45, 46, 47, 48, 49. Node 45 is connected to node 46. Node 46 is connected to node 47. Node 47 is connected to nodes 48, 49, 50, 51. Node 51 is connected to nodes 52, 53, 54. Node 53 is connected to nodes 54, 55. Node 54 is connected to nodes 55, 56, 57. Node 56 is connected to nodes 57, 58, 59. Node 58 is connected to nodes 59, 60. Node 63 is connected to nodes 64, 65, 66. Node 68 is connected to nodes 69, 70, 71. Node 70 is connected to node 71. Node 71 is connected to node 72. Node 73 is connected to nodes 74, 75. Node 77 is connected to nodes 78, 79, 80. Node 78 is connected to nodes 79, 80, 81. Node 79 is connected to nodes 80, 81, 82. Node 80 is connected to nodes 81, 82, 83. Node 81 is connected to nodes 82, 83, 84, 85, 86, 87, 88. Node 83 is connected to nodes 84, 85, 86. Node 87 is connected to nodes 88, 89, 90. Node 91 is connected to nodes 92, 93, 94. Node 106 is connected to nodes 107, 108, 109. Node 111 is connected to nodes 112, 113, 114. </AST> Output: “‘cpp #include <bits/stdc++.h> using namespace std ; void Kmultiples ( int n , int k ) { int a = n ; for ( int i = 1 ; i <= k ; i ++ ) { cout << n << "*" << i << "=" << a << endl ; int j = 0 ; while ( n >= ( 1 << j ) ) { a += n & ( 1 << j ) ; j ++ ;}}} “‘
Figure 12. Example for learning syntactic structure in XLCoST-Instruct.
@@
Figure 13. Example for learning variable dependency in XLCoST-Instruct.
Appendix D Construction of TransCoder-test-X
@@
Figure 14. Original code in TransCoder-test, evaluating specific code implementations.
@@
Figure 15. Enhanced code in TransCoder-test-X, evaluating various equivalent code implementations.
@@
Figure 16. Alignment of return types for parallel data in TransCoder-test.
@@
Figure 17. Inherent errors in the unit test template present in TransCoder-test.
Appendix E Example in TransCoder-test-X
@@
Figure 18. Example of ExeCoder’s evaluation on TransCoder-test-X.

### v3

@@ Appendix F Example in XLCoST-Instruct
Figure 10: Example for learning code semantic in XLCoST-Instruct.
@@
Figure 11: Example for learning functional semantic in XLCoST-Instruct.
@@
Prompt: Translate the given code from python to cpp. The input contains the source code and a Abstract Syntax Tree of the code. The input Code is marked with <Code>and </Code>. Please note that the code entered is a complete program with main fuction. The Abstract Syntax Tree of the code is marked with <AST>and </AST>. Input: <Code> def Kmultiples ( n , k ) : a = n for i in range ( 1 , k + 1 ) : print ( "{}*{}={}".format ( n , i , a ) ) j = 0 while ( n >= ( 1 << j ) ) : a += n & ( 1 << j ) j += 1 </Code> <AST> G describes a graph among nodes 0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, and 115. In this graph: Node 5 represents code def. Node 6 represents code Kmultiples. Node 8 represents code :. Node 13 represents code (. Node 14 represents code n. Node 15 represents code ,. Node 16 represents code k. Node 17 represents code ). Node 20 represents code N. Node 21 represents code =. Node 22 represents code 16. Node 23 represents code K. Node 24 represents code =. Node 25 represents code 7. Node 26 represents code Kmultiples. Node 29 represents code for. Node 30 represents code i. Node 31 represents code in. Node 33 represents code :. Node 35 represents code (. Node 36 represents code N. Node 37 represents code ,. Node 38 represents code K. Node 39 represents code ). Node 40 represents code a. Node 41 represents code =. Node 42 represents code n. Node 43 represents code range. Node 48 represents code (. Node 49 represents code 1. Node 50 represents code ,. Node 52 represents code ). Node 55 represents code while. Node 57 represents code :. Node 59 represents code k. Node 60 represents code +. Node 61 represents code 1. Node 62 represents code print. Node 64 represents code j. Node 65 represents code =. Node 66 represents code 0. Node 67 represents code (. Node 69 represents code ). Node 72 represents code (. Node 74 represents code ). Node 75 represents code n. Node 76 represents code >=. Node 82 represents code (. Node 84 represents code ). Node 85 represents code a. Node 86 represents code +=. Node 88 represents code j. Node 89 represents code +=. Node 90 represents code 1. Node 92 represents code .. Node 93 represents code format. Node 94 represents code (. Node 95 represents code n. Node 96 represents code ,. Node 97 represents code i. Node 98 represents code ,. Node 99 represents code a. Node 100 represents code ). Node 101 represents code 1. Node 102 represents code <<. Node 103 represents code j. Node 104 represents code n. Node 105 represents code &. Node 107 represents code ". Node108representscode{}*{}={}.Node109representscode". Node 110 represents code (. Node 112 represents code ). Node 113 represents code 1. Node 114 represents code <<. Node 115 represents code j. In this graph: Node 0 is connected to nodes 1, 2, 3, 4. Node 1 is connected to nodes 2, 3, 4, 5, 6. Node 2 is connected to node 3. Node 3 is connected to node 4. Node 4 is connected to node 5. Node 7 is connected to nodes 8, 9, 10, 11, 12. Node 9 is connected to nodes 10, 11. Node 10 is connected to nodes 11, 12, 13. Node 11 is connected to nodes 12, 13, 14. Node 12 is connected to nodes 13, 14. Node 18 is connected to node 19. Node 19 is connected to nodes 20, 21, 22, 23, 24, 25. Node 27 is connected to nodes 28, 29, 30, 31, 32. Node 28 is connected to nodes 29, 30, 31. Node 32 is connected to nodes 33, 34. Node 34 is connected to nodes 35, 36, 37. Node 44 is connected to nodes 45, 46, 47, 48, 49. Node 45 is connected to node 46. Node 46 is connected to node 47. Node 47 is connected to nodes 48, 49, 50, 51. Node 51 is connected to nodes 52, 53, 54. Node 53 is connected to nodes 54, 55. Node 54 is connected to nodes 55, 56, 57. Node 56 is connected to nodes 57, 58, 59. Node 58 is connected to nodes 59, 60. Node 63 is connected to nodes 64, 65, 66. Node 68 is connected to nodes 69, 70, 71. Node 70 is connected to node 71. Node 71 is connected to node 72. Node 73 is connected to nodes 74, 75. Node 77 is connected to nodes 78, 79, 80. Node 78 is connected to nodes 79, 80, 81. Node 79 is connected to nodes 80, 81, 82. Node 80 is connected to nodes 81, 82, 83. Node 81 is connected to nodes 82, 83, 84, 85, 86, 87, 88. Node 83 is connected to nodes 84, 85, 86. Node 87 is connected to nodes 88, 89, 90. Node 91 is connected to nodes 92, 93, 94. Node 106 is connected to nodes 107, 108, 109. Node 111 is connected to nodes 112, 113, 114. </AST> Output: “‘cpp #include <bits/stdc++.h> using namespace std ; void Kmultiples ( int n , int k ) { int a = n ; for ( int i = 1 ; i <= k ; i ++ ) { cout << n << "*" << i << "=" << a << endl ; int j = 0 ; while ( n >= ( 1 << j ) ) { a += n & ( 1 << j ) ; j ++ ;}}} “‘
Figure 12: Example for learning syntactic structure in XLCoST-Instruct.
@@
Figure 13: Example for learning variable dependency in XLCoST-Instruct.
Appendix G Construction of TransCoder-test-X
@@
Figure 14: Original code in TransCoder-test, evaluating specific code implementations.
@@
Figure 15: Enhanced code in TransCoder-test-X, evaluating various equivalent code implementations.
@@
Figure 16: Alignment of return types for parallel data in TransCoder-test.
@@
Figure 17: Inherent errors in the unit test template present in TransCoder-test.
Appendix H Example in TransCoder-test-X
@@
Figure 18: Example of ExeCoder’s evaluation on TransCoder-test-X.